Mr. Jones has a 2-stock portfolio with a total value of $500,000. $185,000 is invested in Stock A and the remainder is invested in Stock B. If standard deviation of Stock A is 20.40%, Stock B is 10.30%, and correlation between Stock A and Stock B is 0.20, what would be the expected risk on Mr. Jones’ portfolio (standard deviation of the portfolio return)?

Calcualte with at least 4 decimal places and round your answer to two decimal places. For example, if your answer is $345.6671 round as 345.67 and if your answer is .05718 or 5.7182% round as 5.72.

Group of answer choices

10.89%

11.66%

11.11%

8.93%

11.55%
Answer #1

Weight of Stock A = 185,000 / 500,000 = 0.37

Weight of stock B = (500,000 - 185,000) / 500,000 = 0.63

Portfolio standard deviation = (W12Q12 + W22Q22 + 2 * W1 *W2 * Correlation* Q1 * Q2)1/2

Portfolio standard deviation = (0.3720.20402 + 0.6320.1032 + 2 * 0.37 * 0.63 * 0.2 * 0.2040 * 0.103)1/2

Portfolio standard deviation = (0.1369*0.041616 + 0.3969*0.010609 + 0.001959)1/2

Portfolio standard deviation = (0.005697 + 0.004211 + 0.001959)1/2

Portfolio standard deviation = 0.1089 or 10.89%